# Leveraging Large Language Models (LLMs) to Empower Training-Free Dataset Condensation for Content-Based Recommendation

### v3

@@ III-C User-Level Condensation
III-C1 Interests Extraction and User Encoder
Interests Extraction: Each user $u$ is associated with a click history $h^{(u)}$, based on which we will utilize LLM to extract the set of interests $\mathcal{I}^{(u)}$ for this user. Specifically, we design the guiding prompt in the following format:
@@ III-C1 Interests Extraction and User Encoder
Then, the click history $h^{(u)}$ will be fed into LLM to generate the interests of user $u$, which could be formulated as follows:
$\mathcal{I}^{(u)}=LLM(h^{(u)}),$ (3)
@@
where $\mathcal{N}$ is the set of items and $f_{\theta}$ is the well-trained user encoder [5, 2, 3].
III-C2 Clustering-Based Synthesis
@@
III-D Discussion
TABLE I: Comparison of Computational Complexity.
@@ III-D Discussion
Kernel-based Distill-CF [14] 1 $\tau_{o}$ $\tau_{o}$
Matching-Based DC [11] $\tau_{i}$ $\tau_{o}$ $\tau_{i}\cdot\tau_{o}$
Doscond [10] 1 $\tau_{o}$ $\tau_{o}$
@@
Complexity Comparison. The condensation methods are categorized into two types [28]: matching-based and kernel-based. The first type generates synthetic datasets via gradient matchings between the model trained on small-condensed dataset and those trained on the large-real dataset [11, 10], which typically involve a complex bi-level optimization. In the inner optimization, the models are trained while the datasets are optimized via outer gradients. In contrast, the kernel-based methods reduce the problem into a single-level optimization, which is always free from model training during condensation. If we denote the number of outer-iteration as $\tau_{o}$ and the number of inner-iteration as $\tau_{i}$, we can respectively obtain the computational complexity for each and we compare our proposed method with several representative condensation approaches. As shown in Table I, we can find that our proposed method is the most efficient since it follows a forward paradigm, thus free from iterative optimization on the synthetic dataset.
IV Experiments
In this section, we conduct experiments to evaluate TF-DCon, mainly aiming to answer the following research questions: RQ1: how well the performance can be achieved when trained on datasets condensed by TF-Con? RQ2: how generalizable is the datasets condensed by TF-Con across different recommendation models?
IV-A Experimental Settings
IV-A1 Datasets
To evaluate the performance of our method, we condense the training set of three real-world datasets: MIND [1], Goodreads [29], and MovieLens [30]. The statistics of these datasets are shown in Table II. We split the datasets into $80\%/10\%/10\%$ for training/validation/test. The condensation is mainly conducted on the training sets of datasets and during the testing phase, we still use the users and contents of items from the original datasets.
TABLE II: Dataset statistics. “avg. tok.” denotes the average number of tokens in the content of each item and “avg. his.” denotes the average length of user histories.
@@ IV-A1 Datasets
IV-A2 Baselines
@@
Due to the limited studies on condensation for CBR, we compare our proposed methods with two baselines implemented by ourselves: (i) Random. In the setting of overall performance comparison (Table III), we randomly select a certain portion of users and randomly select a certain portion of tokens to be the contents for each item. In the setting of user-level condensation (Table VII), it is implemented via randomly selecting certain ratios of users while the contents of items remain the same as those in original datasets. In the setting of content-level condensation (Table VI), we implement it by randomly selecting tokens for each item. (ii) Majority. In the overall performance comparison, we implement this method by sampling the users with a large number of interactions and compressing the contents of items by randomly sampling the tokens.
IV-A3 Evaluation Protocol
@@
To evaluate the quality of our method, we test the performance of recommendation models trained on the condensed datasets. In this evaluation, we employ three prevalent content-based recommendation models, i.e., NAML [2], NRMS [5] and Fastformer [3]. They are all designed in a hierarchical manner, i.e., using a content encoder to capture item representations in the history sequence, and a user encoder to fuse the historical item representations into a unified user representation. To be specific, the evaluation is threefold: (1) condense the datasets, (2) train the recommendation models on the condensed datasets, and (3) test the performance of the models. For the evaluation metrics, we follow the common practice [5, 2] to adopt the widely used metrics, i.e., NDCG@$K$ and Recall@$K$, abbreviated as N@$K$ and R@$K$ in the tables, respectively. In this work, we set $K=1,5$ for evaluation on Goodreads and MovieLens, and set $K=5,10$ for MIND. To compare the performance of models trained on condensed datasets and original datasets, we average the percentages of metrics relative to the metrics trained on original datasets. We denote the percentages by “Quality”.
IV-A4 Implementation Details
@@
During training, we employ Adam [31] operator with a learning rate of 5e-3 for three datasets. We select title, abstract, and category feature of each piece of news as the item content for the MIND dataset, and select book or movie name and description feature of each book or movie as the item content for the Goodreads and MovieLens dataset. We concatenate multiple item contents into a single sequence before feeding them to the recommenders. For all three backbone models, we set the embedding dimension of content encoders to 256, the embedding dimension of user encoders to 64, and the negative sampling ratio to 4. We tune the hyperparameters of all base models to attain optimal performance. We average the results of five independent runs for each model. All experiments are conducted on a single NVIDIA GeForce RTX 3090 device.
IV-B Overall Comparison
@@
TABLE III: Overall Performance Comparison on Condensed Datasets. “Quality” denotes the achieved ratio of performance when compared to those trained on original datasets. The detailed condensation ratio could be found in Table IV.
@@
TABLE IV: Comparison between condensed/sampled datasets and original datasets. We use “ORI.”, “RD.” and “MJ.” to represent the statistics of original, random sampled, and majority-selected dataset. “Item” and “User” denote the condensation information of item contents and users, respectively.
@@
To validate the effectiveness of TF-DCon (RQ1), we measure the recommendation performance of CBR models trained in the condensed datasets and their training efficiency. Specifically, we train the recommendation models (i.e., NAML, NRMS, Fastformer) on the condensed datasets and evaluate their performance on the test set of the original datasets. The results are reported in Table III and the detailed condensation ratios are shown in Table IV. Further, we also evaluate the training efficiency on the condensed datasets and the results are summarized in Figure 4.
IV-B1 Overall Performance Comparison
The proposed TF-DCon achieves better performance than the two baselines on three datasets. Notably, as shown in Table III, we approximate over $97\%$ of the original performance with only $5\%$ data on dataset MIND across three different recommender models. From Table IV, we observe that the amazing condensation ratio comes from two parts: reducing the size of contents by $61\%$ (item side) and reducing the size of historical interactions and users (user side) by $98\%$. We observe similar performance on datasets Goodreads and MovieLens. Interestingly, when trained on condensed MovieLens, we may achieve a slightly better performance than those trained on the original dataset. We attribute this to the small size of the original MovieLens dataset since the condensed datasets may serve as a refined and denoised alternative to the original one.
From Table IV, we can find that in each dataset, the size of the user part (i.e., users’ historical interactions) is much larger than the size of the item part (contents of items). Consequently, the predominant size reduction stems from the condensation of the user part. We attribute this promising reduction to the common interests among different users, thus rendering that method of generating representative fake users to replace a variety users reasonably and effective. The condensation ratio of user-side on Goodreads and MovieLens dataset is less promising than the ratio on MIND. This discrepancy is ascribed to the fact that smaller datasets inherently exhibit fewer redundant interests among distinct users, consequently leading to a less pronounced reduction ratio.
IV-B2 Training Efficiency
@@
IV-C Generalization Across Recommender Models
IV-C1 Generalizability of Condensed Datasets
To answer RQ2, we test the performance of CBR models with different architectures trained on the datasets condensed by one CBR model. Specifically, we employ NAML [2] as the backbone model for condensation and test the performance of NAML [2], NRMS [5] and Fastformer [3] trained on the condensed datasets. The results are reported in Table III. Furthermore, we also test the generalizability of content-level and user-level condensation in the ablation study for TF-DCon , as shown in Table VI and Table VII. From the results, we can observe that the datasets condensed by TF-DCon exhibit good generalization on different CBR models. We ascribe this capability to textual data, as texts exclusively encapsulate semantic information while excluding architectural details found in text embeddings.
IV-C2 Versatility of TF-DCon
The proposed TF-DCon is highly flexible since we can adopt different CBR models as the user encoder in the condensation method. We investigate the performances of different models on datasets condensed by TF-DCon with different recommendation models. We utilize NAML, NRMS, and Fastformer as the user encoder for both condensation and evaluation, respectively. We choose dataset MIND to report results in Table V, where we can observe the strong generalization ability of condensed datasets on other models. We can also observe that NAML can outperform the other two models when trained on the most condensed dataset as well as the original dataset.
TABLE V: TF-DCon with different recommender models for condensation. The datasets condensed by different recommender models exhibit similar training performance across different recommender models, verifying the versatility of our method.
Train Set TF-DCon ${}_{NAML}$ TF-DCon ${}_{NRMS}$ TF-DCon ${}_{Fast.}$ Original
@@ IV-C2 Versatility of TF-DCon
IV-D Ablation Study
To examine how user-level condensation and content-level condensation affect the overall performance, we performed an ablation study on them, respectively. We examine the performance of models trained on datasets generated by only user-level condensation or content-level condensation. Specifically, content-level condensation only condense the contents of items and the historical interactions. The results are reported in Table VI, where we observe that only condensing the item contents can also achieve comparable performance. The condensation ratio is relatively not promising and we attribute this to the information capacity of the contents which restricts the size of textual data.
@@ IV-D Ablation Study
For user-level condensation, we only condense the users into fake users and their corresponding interactions while the contents are preserved. As shown in Table VII, we can find that the condensation ratio is astonishingly good while we can still achieve good performance. Especially, on dataset MIND, we can achieve over 97% performance of the original datasets with 98% size reduction of the user-side dataset (i.e., the historical interactions of users ) when tested on Fastformer. When tested on NAML and NRMS, the performance can also exceed over 93% of those trained on the original dataset. Besides, the condensation ratios on Goodreads and MovieLens are also good with a comparable performance. We attribute the promising condensation ratios to similar interests among different users. Therefore, when we synthesize a fake user to represent a group of users with common interests, we can provide adequate information for model training.
@@
TABLE VI: Ablation Study on Content-Level Condensation.
@@
TABLE VII: Ablation Study on User-Level Condensation.
@@
IV-E Further Investigation

### v4

@@ 4.2 User-Level Condensation
4.2.1 Interests Extraction and User Encoder
Interests Extraction: Each user $u$ is associated with a click history $h^{(u)}$, based on which we will utilize ChatGPT to extract the set of interests $\mathcal{I}^{(u)}$ for this user. Specifically, we design the guiding prompt in the following format:
@@ 4.2.1 Interests Extraction and User Encoder
Then, the click history $h^{(u)}$ will be fed into ChatGPT to generate the interests of user $u$, which could be formulated as follows:
$\mathcal{I}^{(u)}=ChatGPT(h^{(u)}),$ (3)
@@
where $\mathcal{N}$ is the set of items and $f_{\theta}$ is the well-trained user encoder (Wu et al., 2019b, a, 2021).
4.2.2 Clustering-Based Synthesis
@@
4.3 Computational Complexity Comparison.
Table 1: Comparison of Computational Complexity.
@@
Kernel-based Distill-CF (Sachdeva et al., 2022) 1 $\tau_{o}$ $\tau_{o}$
Matching-Based DC (Zhao et al., 2021) $\tau_{i}$ $\tau_{o}$ $\tau_{i}\cdot\tau_{o}$
Doscond (Jin et al., 2022a) 1 $\tau_{o}$ $\tau_{o}$
@@
In this section, we discuss about the time complexity of TF-DCon and previous dataset condensation methods. The complexity in Table 1 is calculated by counting the times of generation/training based on the whole dataset. Previous condensation methods are categorized into two types (Cui et al., 2023b): matching-based and kernel-based. The first type generates synthetic datasets via gradient matchings between the model trained on small-condensed dataset and those trained on the large-real dataset (Zhao et al., 2021; Jin et al., 2022a), which typically involve a complex bi-level optimization. In the inner optimization, the models are trained while the datasets are optimized via outer gradients. In contrast, the kernel-based methods reduce the problem into a single-level optimization, which is always free from model training during condensation. If we denote the number of outer-iteration as $\tau_{o}$ and the number of inner-iteration as $\tau_{i}$, we can respectively obtain the computational complexity for each and we compare our proposed method with several representative condensation approaches. As shown in Table 1, we can find that our proposed method is the most efficient since it follows a forward paradigm, thus free from iterative optimization on the synthesized dataset.
5 Experiments
Table 2: Dataset statistics. “avg. tok.” denotes the average number of tokens in the content of each item and “avg. his.” denotes the average length of user histories.
@@ 5 Experiments
In this section, we conduct experiments to evaluate TF-DCon, mainly aiming to answer the following research questions: RQ1: how well the performance can be achieved when trained on datasets condensed by TF-Con? RQ2: how generalizable is the condensed datasets across different recommendation models? RQ3: How efficient is TF-DCon when compared to previous condensation methods? RQ4: Can TF-DCon work well with other less powerful LLM?
5.1 Experimental Settings
5.1.1 Datasets
To evaluate the performance of our method, we condense the training set of three real-world datasets: MIND (Wu et al., 2020), Goodreads (Wan and McAuley, 2018), and MovieLens (Harper and Konstan, 2015). The statistics of these datasets are shown in Table 2. We split the datasets into $80\%/10\%/10\%$ for training/validation/test, respectively. The condensation is mainly conducted on the training sets of datasets and during the testing phase, we still use the users and contents of items from the original datasets.
5.1.2 Baselines
@@
Due to the limited studies on condensation for CBR, we compare our proposed methods with three baselines implemented by ourselves: (i) Random. In the setting of overall performance comparison (Table 3), we randomly select a certain portion of users and randomly select a certain portion of tokens to be the contents for each item. In the setting of user-level condensation (Table 10), it is implemented via randomly selecting certain ratios of users while the contents of items remain the same as those in original datasets. In the setting of content-level condensation (Table 9), we implement it by randomly selecting tokens for each item. (ii) Majority. In the overall performance comparison, we implement this method by sampling the users with a large number of interactions and compressing the contents of items by randomly sampling the tokens. (iii) TextEmbed. For the efficiency comparison, we adapt this text-embedding-based condensation method (Li and Li, 2021), originally designed for classification, by replacing the cross-entropy loss with BPR loss. Due to its limited performance in our context, TextEmbed is included solely in the efficiency analysis and excluded in the main performance comparison. For fair comparison, we maintained an equivalent number of synthesized users as employed in our method, aiming to learn a representation for each user as well as a representation for the corresponding candidate item.
5.1.3 Evaluation Protocol
@@
To evaluate the quality of our method, we test the performance of recommendation models trained on the condensed datasets. In this evaluation, we employ three prevalent content-based recommendation models, i.e., NAML (Wu et al., 2019a), NRMS (Wu et al., 2019b) and Fastformer (Wu et al., 2021). They are all designed in a hierarchical manner, i.e., using a content encoder to capture item representations in the history sequence, and a user encoder to fuse the historical item representations into a unified user representation. To be specific, the evaluation is threefold: (1) condense the datasets, (2) train the recommendation models on the condensed datasets, and (3) test the performance of the models. For the evaluation metrics, we follow the common practice (Wu et al., 2019b, a) to adopt the widely used metrics, i.e., NDCG@$K$ and Recall@$K$, abbreviated as N@$K$ and R@$K$ in the tables, respectively. In this work, we set $K=1,5$ for evaluation on Goodreads and MovieLens, and set $K=5,10$ for MIND. To compare the performance of models trained on condensed datasets and original datasets, we average the percentages of metrics relative to the metrics trained on original datasets. We denote the percentages by “Quality”.
5.1.4 Implementation Details
@@
During dataset condensation, we utilize GPT-3.5 for condensing. During training, we employ Adam (Kingma and Ba, 2015) operator with a learning rate of 5e-3 for three datasets. We select title, abstract, and category feature of each piece of news as the item content for the MIND dataset, and select book or movie name and description feature of each book or movie as the item content for the Goodreads and MovieLens dataset. We concatenate multiple item contents into a single sequence before feeding them to the recommenders. For all three backbone models, we set the embedding dimension of content encoders to 256, the embedding dimension of user encoders to 64, and the negative sampling ratio to 4. We tune the hyperparameters of all base models to attain optimal performance. We average the results of five independent runs for each model. All experiments are conducted on a single NVIDIA GeForce RTX 3090 device.
5.2 Overall Comparison (RQ1)
@@
Table 3: Overall Performance Comparison on Condensed Datasets. “Quality” denotes the achieved ratio of performance when compared to those trained on original datasets. The detailed condensation ratio could be found in Table 4.
@@
Table 4: Comparison between condensed/sampled datasets and original datasets. We use “ORI.”, “RD.” and “MJ.” to represent the statistics of original, random sampled, and majority-selected dataset. “Item” and “User” denote the condensation information of item contents and users, respectively.
@@
To validate the effectiveness of TF-DCon , we measure the recommendation performance of CBR models trained in the condensed datasets and their training efficiency. Specifically, we train the recommendation models (i.e., NAML, NRMS, Fastformer) on the condensed datasets and evaluate their performance on the test set of the original datasets. The results are reported in Table 3 and the detailed condensation ratios are shown in Table 4. Further, we also evaluate the training efficiency on the condensed datasets and the results are summarized in Figure 4.
5.2.1 Overall Performance Comparison
The proposed TF-DCon achieves better performance than the two baselines on three datasets. We utilize “’Quality” to evaluate the quality of condensed dataset, which is computed by dividing each of the four metrics by the original performance, and then averaging them. Notably, as shown in Table 3, we approximate over $97\%$ of the original performance with only $5\%$ data on the MIND dataset across three different recommender models. From Table 4, we observe that the amazing condensation ratio comes from two parts: reducing the size of contents by $61\%$ (item side) and reducing the size of historical interactions and users (user side) by $98\%$. We observe similar performance on datasets Goodreads and MovieLens. Interestingly, when trained on condensed MovieLens, we may achieve a slightly better performance than those trained on the original dataset. We attribute this to the small size of the original MovieLens dataset since the condensed datasets may serve as a refined and denoised alternative to the original one.
From Table 4, we can find that in each dataset, the size of the user part (i.e., users’ historical interactions) is much larger than the size of the item part (contents of items). Consequently, the predominant size reduction stems from the condensation of the user part. We attribute this promising reduction to the common interests among different users, thus rendering that method of generating representative fake users to replace a variety users reasonably and effective. The condensation ratio of user-side on Goodreads and MovieLens dataset is less promising than the ratio on MIND. This discrepancy is ascribed to the fact that smaller datasets inherently exhibit fewer redundant interests among distinct users, consequently leading to a less pronounced reduction ratio. Specifically, within the MIND dataset, $M$ users may share analogous preferences and interaction histories. In contrast, this number diminishes to $N$ within the Goodreads and MovieLens datasets ($M>>N$). Consequently, in the condensed dataset derived from MIND, the synthesis of a single user effectively encapsulates the information equivalent to that of $M$ users in the original dataset. In comparison, the synthesis of one user in the condensed datasets derived from Goodreads and MovieLens covers the information of $N$ users in each dataset, respectively.
5.2.2 Training Efficiency
@@
5.3 Generalization Across Recommender Models (RQ2)
5.3.1 Generalizability of Condensed Datasets
We test the performance of CBR models with different architectures trained on the datasets condensed by one CBR model. Specifically, we employ NAML (Wu et al., 2019a) as the backbone model for condensation and test the performance of NAML (Wu et al., 2019a), NRMS (Wu et al., 2019b) and Fastformer (Wu et al., 2021) trained on the condensed datasets. The results are reported in Table 3. Furthermore, we also test the generalizability of content-level and user-level condensation in the ablation study for TF-DCon , as shown in Table 9 and Table 10. From the results, we can observe that the datasets condensed by TF-DCon exhibit good generalization on different CBR models. We ascribe this capability to textual data, as texts exclusively encapsulate semantic information while excluding architectural details found in text embeddings.
5.3.2 Versatility of TF-DCon
The proposed TF-DCon is highly flexible since we can adopt different CBR models as the user encoder in the condensation method. We investigate the performances of different models on datasets condensed by TF-DCon with different recommendation models. We utilize NAML, NRMS, and Fastformer as the user encoder for both condensation and evaluation, respectively. We choose the MIND dataset to report results in Table 5, where we can observe the strong generalization ability of condensed datasets on other models. We can also observe that NAML can outperform the other two models when trained on the most condensed dataset as well as the original dataset.
Table 5: TF-DCon with different recommender models for condensation. The datasets condensed by different recommender models exhibit similar training performance across different recommender models, verifying the versatility of our method.
Train Set TF-DCon NAML TF-DCon NRMS TF-DCon Fast. Original
@@ 5.3.2 Versatility of TF-DCon
5.4 Condensation Efficiency (RQ3)
Since TF-DCon is a training-free condensation method, we examine its efficiency and compare it with previous condensation method in this section. Specifically, we compare the performance and running time of TF-DCon with TextEmbed (Li and Li, 2021), which follows the most prevalent optimization paradigm (Zhao et al., 2021; Jin et al., 2022b).We report the results on three datasets with NAML as the recommendation model, as shown in Table 7. We can observe that TF-DCon is much faster than the iterative TextEmbed and the performance is also more superior. The efficiency of TF-DCon originates from its training-free design, which is free from the nested optimization involved in TextEmbed.
5.5 Generalization Across Language Models(RQ4)
In this section, we aim to examine:
  • The generalization of TF-DCon with other language models.
  • The impact of evolutionary prompt generations on condensation performance.
Specifically, we equip TF-DCon with Llama-2-7b (Touvron et al., 2023) and utilize different generation of prompts from EvoPro (Algorithm 2) for the content-level condensation.
To examine the performance of TF-DCon with Llama-2 and the influence of EvoPro, we conduct experiments on three datasets with recommendation model NAML. The results are reported in Table 6, where “-evo-0” denotes the method with the initial handcrafted prompt and “-evo-i” denotes the method with $gen_{i}$-prompt. On the one hand, we can observe that the method “TF-DCon” with ChatGPT outperforms all the variants with Llama-2, which is reasonable since the capability of text comprehension and generation of ChatGPT is superior to Llama-2-7b. However, the performance with Llama-2-7b (“-evo-i”) remains close to that achieved with ChatGPT (“TF-DCon”), indicating that our method is effective even with smaller-scale language models. On the other hand, we can observe that the next generation of prompt can consistently outperform its “parent”: “-evo-1” is better than “-evo-0” while “-evo-2” is better than “-evo-1”. This demonstrates that the evolution of the prompts is beneficial in adapting LLM for condensation.
Table 6: TF-DCon with Open-sourced LLM and Evolutionary Prompting for Condensation.
Dataset MIND, $r$=5Goodreads, $r$=27MovieLens, $r$=26%
Metric N@1 N@5 R@1 R@5 N@1 N@5 R@1 R@5 N@1 N@5 R@1 R@5
-evo-0 0.2902 0.3494 0.4278 0.5919 0.5098 0.7903 0.4419 0.9978 0.8330 0.8359 0.1802 0.7312
-evo-1 0.2934 0.3526 0.4315 0.6016 0.5122 0.7951 0.4546 0.9979 0.8378 0.8410 0.1862 0.7382
-evo-2 0.2993 0.3545 0.4329 0.6021 0.5258 0.7981 0.4642 0.9982 0.8423 0.8464 0.1857 0.7416
TF-DCon 0.3071 0.3691 0.4377 0.6150 0.5411 0.8033 0.4704 0.9984 0.8480 0.8494 0.1873 0.7475
Table 7: Efficiency and performance comparison between embedding-based method and TF-DCon.
Dataset MIND Goodreads MovieLens
Methods TextEmbedd TF-DCon TextEmbedd TF-DCon TextEmbedd TF-DCon
N@1 0.2678 450.02min 0.3071 67.46min 0.5013 126.24min 0.5411 6.15min 0.7128 64.58min 0.8484 0.47min
N@5 0.3147 0.3691 0.7794 0.8033 0.7569 0.8494
R@1 0.3815 0.4377 0.4301 0.4704 0.1474 0.1873
R@5 0.5198 0.6150 0.9953 0.9984 0.6884 0.7475
Table 8: Analysis on the number of cluster K.
Dataset MIND Goodreads MovieLens
K 1154 2308 4617 5000 1154 2308 4617 5000 1154 2308 4617 5000
N@1 0.2846 0.2913 0.3071 0.3088 0.5005 0.522 0.5411 0.5467 0.8196 0.8345 0.8484 0.8474
N@5 0.3478 0.356 0.3691 0.3701 0.7864 0.7959 0.8033 0.8017 0.8213 0.8399 0.8494 0.8486
R@1 0.4014 0.4135 0.4377 0.4397 0.4344 0.4541 0.4704 0.4804 0.1779 0.1801 0.1873 0.1893
R@5 0.5686 0.5997 0.615 0.6204 0.9984 0.9983 0.9984 0.9986 0.7304 0.7336 0.7475 0.762
5.6 Ablation Study
To examine how user-level condensation and content-level condensation affect the overall performance, we performed an ablation study on them, respectively. We examine the performance of models trained on datasets generated by only user-level condensation or content-level condensation.
@@ 5.6 Ablation Study
Specifically, content-level condensation only condense the contents of items and the historical interactions. The results are reported in Table 9, where we observe that only condensing the item contents can also achieve comparable performance. The condensation ratio is relatively not promising and we attribute this to the information capacity of the contents which restricts the size of textual data. For user-level condensation, we only condense the users into fake users and their corresponding interactions while the contents are preserved. As shown in Table 10, we can find that the condensation ratio is astonishingly good while we can still achieve good performance. Especially, on the MIND dataset, we can achieve over 97% performance of the original datasets with 98% size reduction of the user-side dataset (i.e., the historical interactions of users ) when tested on Fastformer. When tested on NAML and NRMS, the performance can also exceed over 93% of those trained on the original dataset. Besides, the condensation ratios on Goodreads and MovieLens are also good with a comparable performance. We attribute the promising condensation ratios to similar interests among different users. Therefore, when we synthesize a fake user to represent a group of users with common interests, we can provide adequate information for model training.
Moreover, by comparing the performance of different methods in content-level condensation (Table 9) to the performance in user-level condensation (Table 10) and both-level condensation (Table 3), we can find that TF-DCon consistently maintains superior performance while the baseline methods suffer a large performance drop while reducing the user number. We attribute this to the ability of TF-DCon in preserving the preference information for most users via the user synthesis.
NAMLNRMSFastformer
@@
Table 9: Ablation Study on Content-Level Condensation.
@@
Table 10: Ablation Study on User-Level Condensation.
@@
5.7 Further Investigation
5.7.1 Analysis on hyperparameter $\alpha$.